the weights of 1000 children have average of 45 kilograms and a standard deviation of 5 kilograms, how many children weight 35 kg to 55 kg


1
Expert's answer
2022-03-02T14:57:25-0500

"P(35<X<55)=P(\frac{35-45}{5}<Z<\frac{55-45}{5})=P(-2<Z<2)="

"=P(Z<2)-P(Z<-2)=0.9545."

Number of children: "n=1000*0.9545=955."